Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
The sum of two numbers is 36. Their difference is 4. Determine the two unknown numbers.

Answer by checkley77(12844)

X+Y=36
X-Y=4
X=Y+4
(Y+4)+Y=36
2Y+4=36
2Y=36-4
2Y=32
Y=16 ANSWER.
X-16=4
X=4+16
X=20 ANSWER.
PROOF:
16+20=36
36=36